# On computational complexity and average-case hardness of shallow-depth Boson Sampling

### v2

@@ Appendix C Proof of Theorem 3 (worst-case hardness)
Lemma 4 (Brod [33], revised).
@@
$|\bra{J}\hat{\mathcal{U}}(C_{0})\ket{I}|^{2}=c_{Q}|\bra{0}^{\otimes n}Q\ket{0}^{\otimes n}|^{2},$ (18)
where $c_{Q}=2^{-O(N)}$ is a $Q$-dependent constant which can be efficiently computed, $\hat{\mathcal{U}}(C_{0})$ is a unitary operator corresponding to the circuit $C_{0}$, and each of |$\ket{I}$ and |$\ket{J}$ is an $M$-mode Fock-state composed of $N$ single photon states and vacuum states for the rest modes.
@@
In this proof, we revise the main result in [33], incorporating a more quantitative analysis on the coefficient $c_{Q}$, which later determines the imprecision level allowed for the worst-case hardness of shallow-depth Boson Sampling. To do so, we first go through the main result in [33].
Given an arbitrary poly-sized quantum circuit $Q$ on $n$ qubits, there exists a measurement-based quantum computation (MBQC) scheme using constant depth brickwork graph state composed of maximally $m_{Q}=\text{poly}(n)$ number of qubits and $\Theta(m_{Q})$ number of CZ gates, with an explicit measurement basis sequence for the universal gate set {CX, T, H} to simulate $Q$ [51, 52, 53] (see, e.g., Fig. 4 (a)). Since the measurement outcomes for gate implementation are random, one needs adaptive measurements based on these measurement outcomes to implement $Q$ (which is indeed the essence of MBQC). However, if post-selecting desired measurement outcomes are allowed, one can rather post-select “good” outcomes that do not require any adaptive measurements to implement $Q$.
This brickwork qubit circuit can also be implemented in linear optical systems via the celebrated KLM scheme [54]. In their scheme, they use a dual-rail encoded state (|$\ket{01}$ or |$\ket{10}$ in the Fock basis) as a qubit, a beam splitter as a one qubit gate, and a 2-depth network of beam splitters for a probabilistic CZ gate which also requires additional two single photon ancillas for post-selecting successful CZ operation. Using this KLM scheme, Ref. [33] proposed a 4-depth linear optical circuit $C_{0}$, which implements the brickwork qubit state by post-selecting successful CZ operation (see Fig. 4 (b) or Fig. 3 in Ref. [33] for more details). Here, the required photon number is $N=\Theta(m_{Q})$ to implement the brickwork qubit state composed of $m_{Q}$ number of qubits and $\Theta(m_{Q})$ number of CZ gates, because each of qubit and CZ gate requires a constant number of photons in the KLM scheme. For more clarity, we illustrate in Fig. 4 the brickwork qubit circuit and corresponding linear optical circuit $C_{0}$ proposed by [33]. Also, the mode number is given by $M=\Omega(N)$, which is required for the dual-rail encoding in the KLM scheme. We remark that the mode number $M$ of $C_{0}$ can be arbitrarily enlarged by considering a circuit that “contains” the circuit in [33], which uses some of the modes for dual-rail encoding, leaving the rest of the modes as vacuum and adding only trivial (i.e., identity) gates at these unwanted modes. As the trivial gates leave the unwanted modes separated, this kind of mode number augmentation does not affect the post-selection probabilities.
To sum up, given an arbitrary quantum circuit $Q$, there exists a 4-depth linear optical circuit $C_{0}$ that can simulate the quantum circuit $Q$, by post-selecting successful CZ operation for all probabilistic CZ gates and post-selecting good measurement outcomes (that do not require any adaptive measurements) to implement each gate in the universal gate set {CX, T, H} composing the circuit $Q$. Here, as the explicit measurement basis sequence of brickwork qubit circuit for the universal gate set {CX, T, H} is given, one can find (i) proper coefficients of beam splitters in $C_{0}$ to implement the universal gate set via the KLM scheme and (ii) good measurement outcomes in the Fock basis. Hence, for the given quantum circuit $Q$, we can efficiently construct the constant-depth circuit $C_{0}$ (by means of determining its coefficients) and $M$-mode $N$-single photon Fock-states |$\ket{I}$ and |$\ket{J}$ such that Eq. (18) holds. Now, $c_{Q}$ becomes a post-selection probability to post-select (i) successful CZ operations for all CZ gates and (ii) good measurement outcomes to implement each gate in the universal gate set composing $Q$ via MBQC. In other words, $c_{Q}$ can be expressed as
@@
where $p_{k}$ denotes post-selection probability to implement $k$ gate (e.g., $p_{\text{CZ}}$ is $2/27$ in [55]), and $\Gamma_{k}$ denotes the number of $k$ gate to implement the circuit $Q$. Hence, given the post-selection probability $p_{k}$ for each gate, by counting the number of each gate to implement the circuit $Q$, $c_{Q}$ can be computed efficiently. To obtain the size of $c_{Q}$, note that we set the photon number $N=\Theta(m_{Q})$ to implement $m_{Q}$ qubit brickwork qubit circuit. As this $m_{Q}$ qubit brickwork qubit circuit can implement at most $O(m_{Q})$ number of gates, $Q$ has at most $O(m_{Q})=O(N)$ number of gates in the universal gate set {CX, T, H}. Also, the number of CZ gates is given as $O(N)$, and thus the total number of gates to implement $Q$ is at most $\sum_{k}\Gamma_{k}=O(N)$. Therefore, $c_{Q}$ has its amplitude $c_{Q}=2^{-O(N)}$, since all the post-selection probabilities $p_{k}$ are given as constants. This concludes the proof.
Next, we use the proposition by [42] which states that estimating the output probability of any BQP circuit is #P-hard for a certain additive imprecision.
Lemma 5 (Kondo et al [42]).
It is #$\rm{P}$-hard to compute $|\bra{0}^{\otimes n}Q\ket{0}^{\otimes n}|^{2}$ for an arbitrary given quantum circuit $Q$ within the additive error less than $2^{-2n}$.
@@
There exists a constant-depth circuit $C_{0}$ in Lemma 4 that can be implemented in $\mathcal{B}\mathcal{B}^{*}$ under a proper mode index permutation. In other words, for an arbitrary given poly-sized $n$-qubit quantum circuit $Q$, there exists a circuit $C_{0}$ in $\mathcal{B}\mathcal{B}^{*}$ that satisfies Eq. (18) for $M=\Omega(N)$ and $N={\rm{poly}}(n)$, where $c_{Q}=2^{-O(N)}$, and |$\ket{I}$ and |$\ket{J}$ are given by $M$-mode $N$-single photon Fock-states.
@@
We prove Lemma 6 by showing that the constant-depth linear optical circuit $C_{0}$ in Lemma 4 can be embedded into a $3$-dimensional grid-structured circuit, which can be implemented in the $\mathcal{B}\mathcal{B}^{*}$ architecture under a suitable permutation of mode indices. For interested readers, we provide the full proof of Lemma 6 at the end of this appendix.
Combining the above results, the proof of Theorem 3 is straightforward. First, given an arbitrary quantum circuit $Q$, by Lemma 4, we can efficiently construct a constant-depth linear optical circuit $C_{0}$ and Fock-states |$\ket{I}$ and |$\ket{J}$ such that Eq. (18) holds. Also, as depicted in Lemma 6, we can efficiently embed this linear optical circuit $C_{0}$ in $\mathcal{B}\mathcal{B}^{*}$ by properly allocating mode indices. This mode index allocation permutes |$\ket{I}$ and |$\ket{J}$; let us denote the permuted version of |$\ket{I}$ and |$\ket{J}$ as |$\ket{I^{\prime}}$ and |$\ket{J^{\prime}}$, respectively. Then, we can construct the linear optical circuit $C_{0}$ in $\mathcal{B}\mathcal{B}^{*}$ such that
$c_{Q}^{-1}p_{\bm{s}_{0}}(C_{0})=|\bra{0}^{\otimes n}Q\ket{0}^{\otimes n}|^{2},$ (20)
for collision-free input $\bm{t}$ and output $\bm{s}_{0}$ corresponding to |$\ket{J^{\prime}}$ and |$\ket{I^{\prime}}$, respectively. Note that, the input and output configuration |$\ket{J^{\prime}}$ and |$\ket{I^{\prime}}$ are fixed and invariant under the quantum circuit $Q$, because the architecture of $C_{0}$ does not depend on $Q$ (more precisely, only the coefficients of $C_{0}$ depend on $Q$). As $c_{Q}=2^{-O(N)}$, if we can estimate $p_{\bm{s}_{0}}(C_{0})$ to within additive imprecision $2^{-(2n+O(N))}=2^{-O(N)}$, we can estimate the right-hand side of Eq. (20) to within additive imprecision $2^{-2n}$, which is #P-hard as given in Lemma 5. Therefore, for collision-free input $\bm{t}$ and output $\bm{s}_{0}$, estimating an output probability $p_{\bm{s}_{0}}(C_{0})$ to within additive imprecision $2^{-O(N)}$ for any $C_{0}$ in $\mathcal{B}\mathcal{B}^{*}$ is #P-hard in the worst case.
@@
C.1 Proof of Lemma 6: Encoding constant-depth worst-case circuit in Lemma 4 to the circuit architecture $\mathcal{B}\mathcal{B}^{*}$
@@
In the following, we prove Lemma 6, showing that the constant depth worst-case linear optical circuit $C_{0}$ argued in [33] and Lemma 4 can be encoded in the Kaleidoscope circuit architecture $\mathcal{B}\mathcal{B}^{*}$ by properly allocating mode indices. To summarize, this can be done in two steps: First, we show how we can map the worst-case linear optical circuit $C_{0}$ into a 3-dimensional “grid” linear optical circuit (i.e., a circuit with only nearest neighbor interactions for each dimension, with open boundaries). Next, we show that a single parallel application of any grid-structured linear optical circuit can be implemented in the unit butterfly circuit architecture $\mathcal{B}$ under a mode index permutation. We then combine these arguments, showing that the worst-case circuit $C_{0}$ can be encoded in $\mathcal{B}\mathcal{B}^{*}$.
@@
C.1.1 Mapping worst-case circuit in Lemma 4 to a grid linear optical circuit
We begin by describing how to map the worst-case linear optical circuit $C_{0}$ in Lemma 4 into a 3-dimensional grid-structured linear optical circuit. To do so, we first depict in Fig. 4 (a) the brickwork qubit circuit for the MBQC [51, 52, 53], which can simulate any BQP circuit using post-selection. Here, each vertex (i.e., black dot) represents input $\ket{+}$ state, and each edge represents the CZ gate applied on two qubits. After these CZ operations, each vertex is measured with rotated $\ket{+}$ basis, where the rotation angle is explicitly given to implement the universal gate set via MBQC.
Figure 4: (a) Schematics of the unit brickwork qubit circuit for measurement-based quantum computation [51, 52, 53]. Each vertex (i.e., black dot) represents $\ket{+}$ qubit state, and each edge represents the CZ gate applied on two qubits. After CZ gate operations, each vertex is measured with rotated $\ket{+}$ basis, where the rotation angle is explicitly given for measurement-based quantum computation. (b) Schematics of the 4-depth linear optical circuit proposed by [33] corresponding to the unit brickwork qubit circuit in (a) implemented by the KLM scheme (for more details, see Ref. [33]). Each vertex (i.e., white dot) represents zero or one photon Fock-state, and each edge represents the beam splitter with arbitrarily chosen coefficients. The initial beam splitter layer (red) corresponds to the input qubit (and ancillas), the following two beam splitter layers (yellow and green) correspond to the CZ gate implementation [55], and the last beam splitter layer (blue) corresponds to the qubit rotation required for the measurement-based quantum computation. After the beam splitter operations, each vertex is measured with the Fock-state basis. (c) Schematics of the unit brickwork linear optical circuit in (b) but mapped in 3-dimensional grid linear optical circuit architecture. Here, the beam splitters are applied in parallel along each dimension, where the application sequence is along the direction $x\rightarrow y\rightarrow z\rightarrow x$.
We can also construct this brickwork qubit circuit in linear optical systems under post-selections. As proposed in [33], the brickwork qubit circuit in Fig. 4 (a) can also be implemented in the linear optical circuit by the KLM scheme combined with post-selection, which is indeed the worst-case 4-depth linear optical circuit $C_{0}$ in Lemma 4. We illustrate the brickwork of this 4-depth linear optical circuit in Fig. 4 (b). Here, each vertex (i.e., white dot) represents zero or one photon Fock-state, and each edge represents the beam splitter with appropriately chosen coefficients. More specifically, the initial beam splitter layer (red) corresponds to the input qubit preparation (and ancillas), the following two beam splitter layers (yellow and green) combined with single photon ancillas correspond to the CZ gate operation [55]. The last beam splitter layer (blue) combined with Fock-state measurement corresponds to the rotated $\ket{+}$ measurement in the KLM scheme, which is required to implement the universal gate set via MBQC.
@@ C.1.1 Mapping worst-case circuit in Lemma 4 to a grid linear optical circuit
C.1.2 Encoding a grid linear optical circuit to the butterfly circuit architecture $\mathcal{B}$
@@
Suppose that a mode number is given by $M=2^{n}$ for $n\in\mathbb{Z}^{+}$. For an $M$-mode single parallel $1$-dimensional grid linear optical circuit $C$, there exist an $M$-mode circuit $C^{\prime}$ in $\mathcal{B}$ and an $M$-mode permutation matrix $\bm{P}$ that satisfy $C=\bm{P}C^{\prime}\bm{P}^{T}$
@@
Let us explain why this permutation $\bm{P}$ can make the grid circuit from the circuit in $\mathcal{B}$. First, suppose that the following is true: For given $L\geq 2$ and $j$, each of the sub-circuit of depth $L-1$ with mode indices $\left\llbracket 2^{L-1}(2j-2)+1,2^{L-1}(2j-2)+2^{L-1}\right\rrbracket$ and $\left\llbracket 2^{L-1}(2j-1)+1,2^{L-1}(2j-1)+2^{L-1}\right\rrbracket$ forms a grid circuit with mode number $2^{L-1}$; note that, this is trivial for $L=2$ case. By the structure of $\mathcal{B}$, for given $L\geq 2$ and $j$, initially $(2^{L}(j-1)+2^{L-1})$-th mode and $(2^{L}(j-1)+2^{L})$-th mode are connected by a gate (i.e., a beam splitter) at depth $L$. Then, after the ascending-to-descending permutation $\bm{P}_{L,j}$ for given $L\geq 2$ and $j$, now $(2^{L}(j-1)+2^{L-1})$-th mode and $(2^{L}(j-1)+2^{L-1}+1)$-th mode are connected by the gate. Hence, by setting this gate as a non-trivial gate composing the grid circuit and setting the rest of the gates as trivial gates, we now have a grid circuit with size $2^{L}$ for each $L\geq 2$ and $j$, with mode indices given by $\left\llbracket 2^{L}(j-1)+1,2^{L}(j-1)+2^{L}\right\rrbracket$. Moreover, because we permute equivalently for all $j\in\left[2^{n-L}\right]$ for given $L$, these permutations do not affect the structure of $\mathcal{B}$ for larger depth $L+1,L+2,\dots,n$, such that after the permutations, $(2^{L+1}(j-1)+2^{L})$-th mode and $(2^{L+1}(j-1)+2^{L+1})$-th mode are still connected by a gate at depth $L+1$. Therefore, we can inductively repeat this permutation for ascending $j$ and $L$, thereby obtaining the permutation $\bm{P}=\prod_{L=2}^{n}\prod_{j=1}^{2^{n-L}}\bm{P}_{L,j}$ that can implement the grid circuit of size $2^{n}$.
@@
Therefore, we can construct the circuit $C^{\prime}$ in $\mathcal{B}$ as follows. First, for depth $L=1$ and $j\in[2^{n-1}]$, we set the gate between the modes $2^{L}(j-1)+2^{L-1}$ and $2^{L}(j-1)+2^{L}$ in $\mathcal{B}$ as $G_{s,t}$ for $s=2^{L}(j-1)+2^{L-1}$ and $t=s+1$. Next, for each depth $L\in\llbracket 2,n\rrbracket$ and $j\in[2^{n-L}]$, we set the gate between the modes $12^{L}(j-1)+2^{L-2}+$ 1 and $12^{L}(j-1)+2^{L-1}+2^{L-2}+$ 1 in $\mathcal{B}$ as $G_{s,t}$ for $s=2^{L}(j-1)+2^{L-1}$ and $t=s+1$, and leave trivial for the rest of the gates. Combining the arguments so far, this circuit $C^{\prime}$ permuted by the proper permutation matrix $\bm{P}$ described above finally gives the desired grid circuit $C=\bm{P}C^{\prime}\bm{P}^{T}$ (see, e.g., Fig. 5 (a)).
@@
In the ascending order of the dimension $i\geq 2$, we note that each mode is departed by $2^{\sum_{p=1}^{i-1}n_{p}}$ mode index from its nearest-neighborhood along the $i$-th dimension. By the structure of $\mathcal{B}$, for given $L\in\left\llbracket\sum_{p=1}^{i-1}n_{p}+1,\sum_{p=1}^{i}n_{p}\right\rrbracket$ and $j\in\left[2^{n-L}\right]$, initially $(2^{L}(j-1)+2^{L-1}-2^{\sum_{p=1}^{i-1}n_{p}}+{l})$-th mode and $(2^{L}(j-1)+2^{L}-2^{\sum_{p=1}^{i-1}n_{p}}+{l})$-th mode are connected by gates at $L$ for all ${l}\in\left[2^{\sum_{p=1}^{i-1}n_{p}}\right]$, where ${l}$ characterizes the indices along all the previous dimensions. Hence, now our permutation rule is that for each $L\in\left\llbracket\sum_{p=1}^{i-1}n_{p}+1,\sum_{p=1}^{i}n_{p}\right\rrbracket$ and $j\in\left[2^{n-L}\right]$, we permute mode indices by
[…]

### v3

@@ Appendix C Proof of Theorem 3 (worst-case hardness)
Lemma 5 (Brod [brod2015complexity], revised).
@@
[…]
@@
In this proof, we revise the main result in [brod2015complexity], incorporating a more quantitative analysis on the coefficient $c_{Q}$, which later determines the imprecision level allowed for the worst-case hardness of shallow-depth Boson Sampling. To do so, we first go through the main result in [brod2015complexity].
Given an arbitrary poly-sized quantum circuit $Q$ on $n$ qubits, there exists a measurement-based quantum computation (MBQC) scheme using constant depth brickwork graph state composed of maximally $m_{Q}=\text{poly}(n)$ number of qubits and $\Theta(m_{Q})$ number of CZ gates, with an explicit measurement basis sequence for the universal gate set {CX, T, H} to simulate $Q$ [raussendorf2001one, broadbent2009universal, childs2005unified] (see, e.g., Fig. 4 (a)). Since the measurement outcomes for gate implementation are random, one needs adaptive measurements based on these measurement outcomes to implement $Q$ (which is indeed the essence of MBQC). However, if post-selecting desired measurement outcomes are allowed, one can rather post-select “good" outcomes that do not require any adaptive measurements to implement $Q$.
[…]
@@
where $p_{k}$ denotes post-selection probability to implement $k$ gate (e.g., $p_{\text{CZ}}$ is $2/27$ in [knill2002quantum]), and $\Gamma_{k}$ denotes the number of $k$ gate to implement the circuit $Q$. Hence, given the post-selection probability $p_{k}$ for each gate, by counting the number of each gate to implement the circuit $Q$, $c_{Q}$ can be computed efficiently. To obtain the size of $c_{Q}$, note that we set the photon number $N=\Theta(m_{Q})$ to implement $m_{Q}$ qubit brickwork qubit circuit. As this $m_{Q}$ qubit brickwork qubit circuit can implement at most $O(m_{Q})$ number of gates, $Q$ has at most $O(m_{Q})=O(N)$ number of gates in the universal gate set {CX, T, H}. Also, the number of CZ gates is given as $O(N)$, and thus the total number of gates to implement $Q$ is at most $\sum_{k}\Gamma_{k}=O(N)$. Therefore, $c_{Q}$ has its amplitude $c_{Q}=2^{-O(N)}$, since all the post-selection probabilities $p_{k}$ are given as constants. This concludes the proof.
Next, we use the proposition by [kondo2022quantum] which states that estimating the output probability of any BQP circuit is #P-hard for a certain additive imprecision.
Lemma 6 (Kondo et al [kondo2022quantum]).
[…]
@@
[…]
@@
We prove Lemma 7 by showing that the constant-depth linear optical circuit $C_{0}$ in Lemma 5 can be embedded into a 33-dimensional grid-structured circuit, which can be implemented in the $\mathcal{B}\mathcal{B}^{*}$ architecture under a suitable permutation of mode indices. For interested readers, we provide the full proof of Lemma 7 in Appendix D.
Combining the above results, the proof of Theorem 3 is now straightforward.
Proof of Theorem 3.
[…]
Remark 2.
In principle, encoding the constant-depth MBQC circuit from Lemma 5 to the logarithmic-depth $\mathcal{BB}^{*}$ circuit, as done in Lemma 7, is not essential for deriving the hardness argument in shallow-depth regimes; one could directly use the constant-depth circuit from Lemma 5 itself to obtain worst-case hardness in such regimes. Nevertheless, since we employ the architecture $\mathcal{BB}^{*}$ to implement required permutations during the worst-to-average-case reduction, we derive worst-case hardness within $\mathcal{BB}^{*}$ to simplify the overall circuit architecture to $(\mathcal{BB}^{*})^{q}$ in Definition 3.
@@
Appendix D Proof of Lemma 7: Encoding constant-depth worst-case circuit in Lemma 5 to the circuit architecture $\mathcal{B}\mathcal{B}^{*}$
@@
In the following, we prove Lemma 7, showing that the constant depth worst-case linear optical circuit $C_{0}$ argued in [brod2015complexity] and Lemma 5 can be encoded in the Kaleidoscope circuit architecture $\mathcal{B}\mathcal{B}^{*}$ by properly allocating mode indices. To summarize, this can be done in two steps: First, we show how we can map the worst-case linear optical circuit $C_{0}$ into a 3-dimensional “grid" linear optical circuit (i.e., a circuit with only nearest neighbor interactions for each dimension, with open boundaries). Next, we show that a single parallel application of any grid-structured linear optical circuit can be implemented in the unit butterfly circuit architecture $\mathcal{B}$ under a mode index permutation. We then combine these arguments, showing that the worst-case circuit $C_{0}$ can be encoded in $\mathcal{B}\mathcal{B}^{*}$.
@@
D.1 Mapping worst-case circuit in Lemma 5 to a grid linear optical circuit
[…]
@@ D.1 Mapping worst-case circuit in Lemma 5 to a grid linear optical circuit
D.2 Encoding a grid linear optical circuit to the butterfly circuit architecture $\mathcal{B}$
@@
Suppose that a mode number is given by $M=2^{n}$ for $n\in\mathbb{Z}^{+}$. For an $M$-mode single parallel 11-dimensional grid linear optical circuit $C$, there exist an $M$-mode circuit $C^{\prime}$ in $\mathcal{B}$ and an $M$-mode permutation matrix $\bm{P}$ that satisfy $C=\bm{P}C^{\prime}\bm{P}^{T}$
@@
Let us explain why this permutation $\bm{P}$ can make the grid circuit from the circuit in $\mathcal{B}$. First, suppose that the following is true: For given $L\geq 2$ and $j$, each of the sub-circuit of depth L1L-1 with mode indices $\left\llbracket 2^{L-1}(2j-2)+1,2^{L-1}(2j-2)+2^{L-1}\right\rrbracket$ and $\left\llbracket 2^{L-1}(2j-1)+1,2^{L-1}(2j-1)+2^{L-1}\right\rrbracket$ forms a grid circuit with mode number $2^{L-1}$; note that, this is trivial for $L=2$ case. By the structure of $\mathcal{B}$, for given $L\geq 2$ and $j$, initially $(2^{L}(j-1)+2^{L-1})$-th mode and $(2^{L}(j-1)+2^{L})$-th mode are connected by a gate (i.e., a beam splitter) at depth $L$. Then, after the ascending-to-descending permutation $\bm{P}_{L,j}$ for given $L\geq 2$ and $j$, now $(2^{L}(j-1)+2^{L-1})$-th mode and $(2^{L}(j-1)+2^{L-1}+1)$-th mode are connected by the gate. Hence, by setting this gate as a non-trivial gate composing the grid circuit and setting the rest of the gates as trivial gates, we now have a grid circuit with size $2^{L}$ for each $L\geq 2$ and $j$, with mode indices given by $\left\llbracket 2^{L}(j-1)+1,2^{L}(j-1)+2^{L}\right\rrbracket$. Moreover, because we permute equivalently for all $j\in\left[2^{n-L}\right]$ for given $L$, these permutations do not affect the structure of $\mathcal{B}$ for larger depth $L+1,L+2,\dots,n$, such that after the permutations, $(2^{L+1}(j-1)+2^{L})$-th mode and $(2^{L+1}(j-1)+2^{L+1})$-th mode are still connected by a gate at depth $L+1$. Therefore, we can inductively repeat this permutation for ascending $j$ and $L$, thereby obtaining the permutation $\bm{P}=\prod_{L=2}^{n}\prod_{j=1}^{2^{n-L}}\bm{P}_{L,j}$ that can implement the grid circuit of size $2^{n}$.
@@
Therefore, we can construct the circuit $C^{\prime}$ in $\mathcal{B}$ as follows. First, for depth $L=1$ and $j\in[2^{n-1}]$, we set the gate between the modes $2^{L}(j-1)+2^{L-1}$ and $2^{L}(j-1)+2^{L}$ in $\mathcal{B}$ as $G_{s,t}$ for $s=2^{L}(j-1)+2^{L-1}$ and $t=s+1$. Next, for each depth $L\in\llbracket 2,n\rrbracket$ and $j\in[2^{n-L}]$, we set the gate between the modes $2^{L}(j-1)+2^{L-2}+1$ and $2^{L}(j-1)+2^{L-1}+2^{L-2}+1$ in $\mathcal{B}$ as $G_{s,t}$ for $s=2^{L}(j-1)+2^{L-1}$ and $t=s+1$, and leave trivial for the rest of the gates. Combining the arguments so far, this circuit $C^{\prime}$ permuted by the proper permutation matrix $\bm{P}$ described above finally gives the desired grid circuit $C=\bm{P}C^{\prime}\bm{P}^{T}$ (see, e.g., Fig. 5 (a)).
@@
In the ascending order of the dimension $i\geq 2$, we note that each mode is departed by $2^{\sum_{p=1}^{i-1}n_{p}}$ mode index from its nearest-neighborhood along the $i$-th dimension. By the structure of $\mathcal{B}$, for given $L\in\left\llbracket\sum_{p=1}^{i-1}n_{p}+1,\sum_{p=1}^{i}n_{p}\right\rrbracket$ and $j\in\left[2^{n-L}\right]$, initially $(2^{L}(j-1)+2^{L-1}-2^{\sum_{p=1}^{i-1}n_{p}}+{l})$-th mode and $(2^{L}(j-1)+2^{L}-2^{\sum_{p=1}^{i-1}n_{p}}+{l})$-th mode are connected by gates at $L$ for all ${l}\in\left[2^{\sum_{p=1}^{i-1}n_{p}}\right]$, where l{l} characterizes the indices along all the previous dimensions. Hence, now our permutation rule is that for each $L\in\left\llbracket\sum_{p=1}^{i-1}n_{p}+1,\sum_{p=1}^{i}n_{p}\right\rrbracket$ and $j\in\left[2^{n-L}\right]$, we permute mode indices by
$2^{L}(j-1)+2^{L-1}+(k-1)2^{\sum_{p=1}^{i-1}n_{p}}+{l}\;\rightarrow\;2^{L}(j-1)+2^{L}-k2^{\sum_{p=1}^{i-1}n_{p}}+{l},$ (23)